# LLMs may Dominate Information Access: Neural Retrievers are Biased Towards LLM-Generated Texts

### v2

@@ 2.3. Statistics and Quality Validation of Datasets
Take the Llama2-generated data as an example, we conduct the statistics and quality validation of the constructed datasets. The analysis of ChatGPT-generated datasets is provided in Appendix A.4, which shows similar observations and conclusions.
@@ 2.3.1. Term-based Statistics and Analysis
We first analyze the term-based similarity between the LLM-generated corpus and the human-written corpus. Specifically, we compute the Jaccard similarity ($\frac{|d^{G}\bigcap d^{H}|}{|d^{G}\bigcup d^{H}|}$) and the overlap ($\frac{|d^{G}\bigcap d^{H}|}{|d^{H}|}$) between each LLM-generated document and orginal human-written document. The distribution of these two term-based similarity metrics is shown in Figure 3. As we can see, both the Jaccard similarity and overlap distributions exhibit normal distribution, with peaks at about 0.4 and 0.6 for NQ320K+AIGC, and about 0.6 and 0.8 for SciFact+AIGC, respectively. These observations suggest that while there is a considerable overlap of terms between the LLM-generated text and the original human-written text, there are also distinct differences, especially noticeable in the NQ320K+AIGC dataset.
@@ 2.3.2. Semantic-based Statistics and Analysis
For the LLM-generated texts, a pivotal consideration is whether they faithfully preserve the underlying semantics of the corresponding human-written corpus. If they indeed do so, we can confidently assign them the same relevance labels as the labels of their corresponding original human-written texts given each query.
To assess this, we first leverage the OpenAI embedding model333text-embedding-ada-002:https://platform.openai.com/docs/api-reference/embeddings to acquire semantic embeddings for both the LLM-generated and human-written corpora. Subsequently, we visualize these embeddings through T-SNE (Van der Maaten and Hinton, 2008) in Figure 4. We observe a strikingly close overlap between the Llama2-generated corpus with the human-written corpus in the latent space. This observation strongly suggests that these Llama2-generated corpora adeptly preserve the original semantics. Moreover, we delve into the cosine similarity of semantic embeddings between the LLM-generated text and their corresponding human-written counterparts. The results, as shown in Figure 5, also indicate a high degree of similarity, with most values exceeding 0.95, affirming the faithful preservation of semantics in LLM-generated text. Hence, for each query-document pair $(q,d^{G})$, we can confidently assign the relevant label $r$ to be the same as that of $(q,d^{H})$.
To further validate the accuracy of the relevance label assignments, we conduct an evaluation of retrieval models on the human-written corpus and the LLM-generated corpus, respectively. The results are presented in Table 2. It is evident that all retrieval models exhibit no significant performance discrepancies in terms of various ranking metrics between the human-written and LLM-generated corpora across all datasets. This observation reinforces the confidence in the quality of our newly constructed datasets.
@@
3. RQ2: Explore The Source Bias
In this section, we conduct extensive experiments on the constructed datasets to explore the source bias from various aspects. With the constructed simulated environment, we first introduce the evaluation metrics to quantify the severity of source bias. We then conduct experiments with different retrieval models on both the first-stage retrieval and the second-stage re-ranking. Finally, we also conduct experiments on corpus generated by LLM with different temperatures to further investigate the source bias.
@@ 3.1. Evaluation Metrics for Source Bias
To quantitatively explore source bias, we calculate the ranking metrics, targeting separately either human-written or LLM-generated corpus. Specifically, for each query, an IR model produces a ranking list that comprises documents from both human-written and LLM-generated corpora. We then calculate ranking metrics separately for human-written and LLM-generated texts, depending on the target data source. When computing metrics for one target (e.g., human-written corpus $\mathcal{C}^{H}$), the data corresponding to the other side (e.g., LLM-generated corpus $\mathcal{C}^{G}$) is masked or disregarded (i.e., mask all the positive label $r\in\mathcal{R}^{G}$ as negative), but the rank of each document is still based on the original ranking list that incorporates a mixture of both types of text. This approach allows us to independently assess the performance of IR models on different corpus sources. For the ranking metrics, we report the top-$K$ Normalized Discounted Cumulative Gain (NDCG@$K$) and Mean Average Precision (MAP@$K$) with $K\in\{1,3,5\}$. For a better understanding of the calculation of the evaluation metrics, we provide a toy example in Appendix A.1.
@@
where the Metric can be NDCG@$K$ and MAP@$K$. Note that Relative $\Delta>0$ means retrieval models rank human-written texts higher, and Relative $\Delta<0$ indicates LLM-generated texts are ranked higher. The greater the absolute value of Relative $\Delta$, the greater the ranking performance difference on two corpora.
3.2. Bias in Different Retrieval Models
Our initial set of experiments is aimed at conducting a comprehensive analysis of different retrieval models within the context of the LLM era. These models encompass both traditional lexical approaches and modern neural IR models. The results targeted on either human-written or LLM-generated corpus on the SciFact+AIGC and NQ320K+AIGC are reported in Table 3. Based on these results, we have the following noteworthy observations and conclusions:
Lexical approaches tend to favor human-written documents. When the retrieved corpora contain both human-written and LLM-generated texts, TF-IDF and BM25 achieve better results on the human-written corpus. This phenomenon is observed in terms of almost all ranking metrics across both the SciFact+AIGC and NQ320K+AIGC datasets. A plausible explanation for this phenomenon lies in the term-based distinctions between text generated by LLMs and human-written content, as evident in Figure 3. On the other hand, the queries are crafted by humans and thus exhibit a style more closely aligned with human-written text.
Neural retrieval models are biased towards LLM-generated texts. An intriguing phenomenon emerges when evaluating neural retrieval models that rely on semantic matching. There is a significant performance gap between their performance on the human-written corpus and LLM-generated corpus. Note that all neural retrieval models achieve a relative performance gain of over $30\%$ when targeting LLM-generated corpus as compared to human-written corpus. These findings suggest an inherent bias in neural retrieval models towards LLM-generated text, which we named the source bias. This bias may arise from the potential exploitation of specific patterns or ‘semantic shortcuts’ embedded in LLM-generated text, given that both neural retrieval models (based on PLMs) and LLMs share a similar Transformer-based (Vaswani et al., 2017) architecture and a similar pretraining paradigm. Moreover, it seems that the information in texts is compressed at the semantic level during LLM generation, resulting in a clearer understanding by neural retrieval models. We will delve deeper into the cause of this bias in the subsequent section.
This set of experiments provides valuable insights into the performance of different retrieval models in the LLM era, underscoring the presence of the source bias within neural retrieval models. Next, we will further uncover more scenarios that exhibit this bias.
Table 4. Performance comparison of different neural retrieval models for mixed human-written and ChatGPT-generated corpora on SciFact+AIGC dataset. The annotations in the table are consistent with Table 3.
@@ 3.2. Bias in Different Retrieval Models
Table 5. Bias evaluation of re-ranking models on mixed human-written and Llama2-generated corpora of SciFact+AIGC dataset. The re-ranking methods rerank the top-100 retrieved hits from a first-stage BM25 model. The annotations in the table are consistent with Table 3.
Metrics Target Corpus Base Re-ranking
BM25 +MiniLM +monoT5
NDCG@1 Human-Written 26.7 21.3 19.7
LLM-Generated 21.0 32.7 39.7
$\text{Relative}~{}~{}\Delta$ 23.9 -42.2 -67.3
NDCG@3 Human-Written 40.3 42.8 45.9
LLM-Generated 38.8 47.8 52.9
$\text{Relative}~{}~{}\Delta$ 3.8 -11.0 -14.2
NDCG@5 Human-Written 44.4 46.9 49.0
LLM-Generated 41.5 50.2 54.7
$\text{Relative}~{}~{}\Delta$ 6.8 -6.8 -11.0
MAP@1 Human-Written 25.7 20.8 18.9
LLM-Generated 19.6 30.8 37.8
$\text{Relative}~{}~{}\Delta$ 26.9 -38.8 -66.7
MAP@3 Human-Written 36.7 37.5 39.7
LLM-Generated 34.3 43.6 48.9
$\text{Relative}~{}~{}\Delta$ 6.8 -15.0 -20.8
MAP@5 Human-Written 39.1 40.0 41.6
LLM-Generated 35.9 45.0 50.1
$\text{Relative}~{}~{}\Delta$ 8.5 -11.8 -18.5
3.3. Bias in ChatGPT-Generated Corpus
@@
To strengthen our argument that the source bias is not limited to any specific LLM-generated corpus, we extend our investigation to include ChatGPT, another widely adopted and nearly state-of-the-art LLM. We employ ChatGPT to generate a corpus using the same prompts as those utilized with Llama2 in the above experiments. In Appendix A.4, we present the detailed data quality evaluation of the ChatGPT-generated corpus. Subsequently, in Table 4, we report the evaluation results on the SciFact+AIGC dataset, which contains both human-written and ChatGPT-generated texts. Once again, the results clearly indicate a bias within neural retrieval models, favoring the corpus generated by ChatGPT across all ranking metrics. This observation provides additional substantiation of the presence of source bias within these neural retrieval models.
@@
Table 6. Bias evaluation of re-ranking models on mixed human-written and ChatGPT-generated corpora of SciFact+AIGC dataset. The re-ranking methods rerank the top-100 retrieved hits from a first-stage BM25 model. The annotations in the table are consistent with Table 3.
Metrics Target Corpus Base Re-ranking
BM25 +MiniLM +monoT5
NDCG@1 Human-Written 24.3 18.3 21.3
LLM-Generated 24.3 35.7 39.3
$\text{Relative}~{}~{}\Delta$ 0.0 -64.4 -59.4
NDCG@3 Human-Written 38.5 41.4 46.4
LLM-Generated 40.2 50.1 54.2
$\text{Relative}~{}~{}\Delta$4.3 -19.0 -15.5
NDCG@5 Human-Written 42.7 45.6 48.9
LLM-Generated 42.7 53.0 56.1
$\text{Relative}~{}~{}\Delta$ 0.0 -15.0 -13.7
MAP@1 Human-Written 23.7 17.9 20.5
LLM-Generated 23.1 33.8 37.8
$\text{Relative}~{}~{}\Delta$ 2.6 -61.5 -59.3
MAP@3 Human-Written 34.8 35.8 40.3
LLM-Generated 35.8 45.9 50.0
$\text{Relative}~{}~{}\Delta$2.8 -24.7 -21.5
MAP@5 Human-Written 37.3 38.3 41.7
LLM-Generated 37.3 47.6 51.4
$\text{Relative}~{}~{}\Delta$ 0.0 -21.7 -20.8
@@ 3.4. Bias in Re-Ranking Stage
In a typical IR system, there are two primary stages of document filtering. The first stage involves a retriever, responsible for document recall, while the second stage employs a re-ranker, which fine-tunes the ordering of documents within the initially retrieved set. Our investigation has revealed the presence of the source bias in the initial document filtering process (i.e., retrieval stage). Hence, a natural pivotal research question arises: does this bias extend to the ranking models used in the second-stage document filtering process (re-ranking stage), within IR systems?
To delve into this, we choose two representative re-ranking models: MiniLM (Wang et al., 2020) and monoT5 (Nogueira et al., 2020) to rerank the top-100 document list retrieved by a first-stage BM25 model. MiniLM is a cross-encoder model pre-trained on MS MARCO (Nguyen et al., 2016) using a knowledge distillation on three teacher models: BERT-base (Devlin et al., 2019), BERT-large (Devlin et al., 2019), and ALBERT-large (Lan et al., 2019) models. monoT5 is a sequence-to-sequence re-ranker that uses T5 (Raffel et al., 2020) to calculate the relevance score, which is also pretrained on MS MARCO. The results on the SciFact+AIGC dataset with Llama-generated corpus and ChatGPT-generated corpus are presented in Table 5 and Table 6, respectively.
From the results, while even the first-stage retrievers (BM25) may exhibit a preference for human-written content, the second-stage re-rankers once again demonstrate a bias in favor of LLM-generated content. Remarkably, the bias in re-ranking models appears to be more pronounced, as evidenced by the relative percentage difference of $67.3\%$ and $59.4\%$ in NDCG@$1$ for monoT5 in Table 5 and Table 6, respectively. These findings further solidify the pervasiveness of source bias in neural ranking models that rely on PLMs, regardless of the retrieval stage or re-ranking stage.
(a) Relative $\Delta$ on NDCG@$1$
(b) Relative $\Delta$ on MAP@$1$
Figure 6. Performance comparison (NDCG@$1$ and MAP@$1$) of neural retrieval models on SciFact+AIGC dataset under different temperatures of Llama2-generated corpus. A larger negative the Relative $\Delta$ indicates a more severe source bias.
3.5. Bias in Different Temperatures
Temperature is an important hyperparameter used to control the randomness and creativity of text generation in LLMs. A lower temperature sharpens the probability distribution of predicted words, favoring the most likely words. Conversely, a higher temperature flattens the probability distribution, allowing all words to have a more equal chance of being selected. To investigate the impact of temperature on the source bias, we conduct experiments in which Llama2 generated different corpora with varying temperatures (i.e., $T\in\{0.2,1.0\}$). We then examine different neural retrieval models by combining the Llama2-generated corpus with the human-written corpus one by one. The relative gaps in NDCG@$1$ and MAP@$1$ on the SciFact+AIGC dataset are presented in Figure 6. In Appendix A.4, we present detailed data statistics and quality evaluation of the Llama2-generated corpus ($T=1.0$). We also provide the raw results of these retrieval models for the mixed human-written and Llama2-generated corpora ($T=1.0$) in  Table 9 of Appendix  A.5.
The results suggest that all neural retrieval models still exhibit a bias towards LLM-generated text, regardless of the temperature. Furthermore, there is a reduced bias on most neural retrieval models when applied to LLM-generated text with higher temperatures, especially on the re-ranking models. Note that if the temperature coefficient is higher ($T>1.0$), the LLM will not be able to generate coherent meaningful texts, and the generated results will be very random. These findings provide us with potential starting points for analyzing this bias and we will illustrate this in the next section.

### v3

@@ 2.3. Statistics and Quality Validation of Datasets
Take the Llama2-generated data as an example, we conduct the statistics and quality validation of the constructed datasets. The analysis of ChatGPT-generated datasets shows similar observations and conclusions and is omitted due to the page limitation.
@@ 2.3.1. Term-based Statistics and Analysis
We first analyze the term-based similarity between the LLM-generated corpus and the human-written corpus. Specifically, we compute the Jaccard similarity ($\frac{|d^{G}\bigcap d^{H}|}{|d^{G}\bigcup d^{H}|}$) and the overlap ($\frac{|d^{G}\bigcap d^{H}|}{|d^{H}|}$) between each LLM-generated document and orginal human-written document. As shown in Figure 3, both the Jaccard similarity and overlap distributions exhibit normal distribution, with peaks at about 0.6 and 0.8 for SciFact+AIGC, and about 0.4 and 0.6 for NQ320K+AIGC, respectively. These observations suggest that while there is a considerable overlap of terms between the LLM-generated text and the original human-written text, there are also distinct differences, especially noticeable in the NQ320K+AIGC dataset.
(a) SciFact+AIGC
(b) NQ320K+AIGC
@@ 2.3.2. Semantic-based Statistics and Analysis
For the LLM-generated texts, a pivotal consideration is whether they faithfully preserve the underlying semantics of the corresponding human-written corpus. If they indeed do so, we then can confidently assign them the same relevancy labels as the labels of their corresponding original human-written texts given each query.
To assess this, we first leverage the OpenAI embedding model333text-embedding-ada-002:https://platform.openai.com/docs/guides/embeddings to acquire semantic embeddings for both the LLM-generated and human-written texts. Subsequently, we visualize these embeddings through T-SNE (Van der Maaten and Hinton, 2008) in Figure 4. We observe a strikingly close overlap between the Llama2-generated corpus and the human-written corpus in the latent space. This observation strongly suggests that these LLM-generated corpora adeptly preserve the original semantics. Moreover, we delve into the cosine similarity of semantic embeddings between the LLM-generated text and their corresponding human-written counterparts. The results, as shown in Figure 5, also indicate a high degree of similarity, with most values exceeding 0.95, affirming the faithful preservation of semantics in LLM-generated text. Hence, for each query-document pair $(q,d^{G})$, we can confidently assign the relevancy label $r$ to be the same as that of $(q,d^{H})$.
Table 3. Verification of semantics and text quality with human evaluation. The numbers in parentheses represent the proportion agreed upon by all three human annotators for each option.
SciFact+AIGC NQ320K+AIGC
Which document is more relevant to the given query?
Human LLM Equal Human LLM Equal
$0.0\%(0.0\%)$ $0.0\%(0.0\%)$ $100.0\%(82.0\%)$ $2.0\%(0.0\%)$ $0.0\%(0.0\%)$ $98.0\%(81.6\%)$
Which document exhibits higher quality by considering the following aspects:
linguistic fluency, logical coherence, and information density?
Human LLM Equal Human LLM Equal
$8.0\%(0.0\%)$ $6.0\%(0.0\%)$ $86.0\%(46.5\%)$ $4.0\%(0.0\%)$ $6.0\%(0.0\%)$ 90.0%(60.%)90.$0$ ( 60 . % )
@@
2.3.3. Retrieval Performance Evaluation
To further validate the accuracy of the relevancy label assignments, we conduct an evaluation of retrieval models on the human-written corpus and the LLM-generated corpus, respectively. The following representative retrieval models are adopted in the experiments: (1) Lexical Retrieval Models: TF-IDF (Sparck Jones, 1972) and BM25 (Robertson et al., 2009) and (2) Neural Retrieval Models: ANCE (Xiong et al., 2020), BERM (Xu et al., 2023), TAS-B (Hofstätter et al., 2021), Contriever (Izacard et al., 2021).
The results on each sole source corpus on the proposed two new benchmarks are presented in Table 2. It is evident that all retrieval models exhibit no significant performance discrepancies in terms of various ranking metrics between the human-written and LLM-generated corpora across all datasets. This observation reinforces the confidence in the quality of our newly constructed datasets.
2.3.4. Human Evaluation
Note that in our constructed datasets, LLMs were instructed to rewrite human-written texts based solely on the original human-written text, without any query-related input, thereby preventing the additional query-specific information during rewriting. Moreover, to further verify this, we conduct a human evaluation. Specifically, we randomly select 50 <query, human-written document, LLM-generated document> triples from each dataset. The human annotators, comprising the authors and their highly educated colleagues, are asked to determine which document is more semantically relevant to the given query. The options are “Human”, “LLM”, or “Equal”. During the evaluation, annotators are unaware of the source of each document. Each triple is labeled at least by three different annotators, with the majority vote determining the final label. The results in Table 3, confirm that both sources of texts have almost the same semantic relevance to the given queries, which guarantees the fairness of our following exploration of source bias.
Additionally, we also conduct further human evaluations specifically focused on text quality. The human annotators are asked to determine “Which document exhibits higher quality by considering the following aspects: linguistic fluency, logical coherence, and information density?” The notation process is the same as above, and the results are summarized in Table 3. The results indicate no significant distinction between LLM-generated and human-written content on text quality, demonstrating consistency across both sources. In fact, we also analyze the data cases and find that LLMs typically alter only parts of the vocabulary, leading to minor stylistic differences without impacting the core content, which can be further verified with these human evaluations.
3. RQ2: Uncovering Source Bias
In this section, we conduct extensive experiments on the constructed datasets to explore the source bias from various aspects. With the constructed simulated environment, we first introduce the evaluation metrics to quantify the severity of source bias. We then conduct experiments with different retrieval models on both the first-stage retrieval and the second-stage re-ranking.
@@ 3.1. Evaluation Metrics for Source Bias
To quantitatively explore source bias, we calculate ranking metrics, targeting separately either human-written or LLM-generated corpus. Specifically, for each query, an IR model produces a ranking list that comprises documents from mixed corpora. We then calculate top-$K$ Normalized Discounted Cumulative Gain (NDCG@$K$) and Mean Average Precision (MAP@$K$), for $K\in\{1,3,5\}$, independently for each corpus source. When assessing one corpus (e.g., human-written), documents from the other (e.g., LLM-generated) are treated as non-relevant, though the original mixed-source ranking order is maintained. This approach allows us to independently assess the performance of IR models on each corpus source.
@@
where the Metric can be NDCG@$K$ and MAP@$K$. Note that Relative $\Delta>0$ means retrieval models rank human-written texts higher, and Relative $\Delta<0$ indicates LLM-generated texts are ranked higher. The greater the absolute value of Relative $\Delta$, the greater the ranking performance difference between two sourced content.
3.2. Bias in Neural Retrieval Models
Table 5. Performance comparison of different neural retrieval models for mixed human-written and ChatGPT-generated corpora on SciFact+AIGC dataset.
@@ 3.2. Bias in Neural Retrieval Models
In our assessment of various retrieval models on SciFact+AIGC and NQ320K+AIGC datasets, we observe distinct phenomena when evaluating against human-written and LLM-generated corpora, as reported in Table 4. Our key findings are as follows:
Lexical models prefer human-written texts. Lexical models like TF-IDF and BM25 show a tendency to favor human-written texts over LLM-generated texts across most ranking metrics in both datasets. A plausible explanation for this phenomenon lies in the term-based distinctions between text generated by LLMs and human-written content, as evident in Figure 3. Additionally, the queries are crafted by humans and thus exhibit a style more closely aligned with human-written text.
Neural retrievers are biased towards LLM-generated texts. Neural models, which rely on semantic matching with PLMs, demonstrate a pronounced preference for LLM-generated texts, often performing over 30% better on these compared to human-written texts. These findings suggest an inherent bias in neural retrievers towards LLM-generated text, which we named the source bias. This source bias may stem from PLMs-based neural retrievers and LLMs sharing similar Transformer-based architectures (Vaswani et al., 2017) and pretraining approaches, leading to potential exploitation of semantic shortcuts in LLM-generated text during semantic matching. Additionally, LLMs seem to semantically compress information in a manner that makes it more comprehensible to neural models. A deeper exploration into the causes of source bias is presented in the following section.
@@
To strengthen our conclusion that source bias is not limited to any specific LLM, we extend our investigation to include ChatGPT, another widely adopted and nearly state-of-the-art LLM. We employ ChatGPT to generate a corpus using the same prompts as those utilized with Llama2 in the above experiments. Subsequently, in Table 5, we report the evaluation results on the SciFact+AIGC dataset, which contains both human-written and ChatGPT-generated texts. Once again, the results clearly indicate a bias within neural retrieval models, favoring the corpus generated by ChatGPT across all ranking metrics. This observation provides additional substantiation of the presence of source bias within these neural retrieval models.
Furthermore, we also explore the popular InstructGPT-prompts GitHub Repository, which includes several common prompts for rephrasing passages  [^4] https://github.com/kevinamiri/Instructgpt-prompts. The experimental results in Appendix B show varying degrees of source bias, indicating that common prompts can easily trigger source bias with LLM-generated content. These findings highlight the notable presence of source bias in neural retrieval models towards LLM-generated content.
@@
Table 6. Bias evaluation of re-ranking models on SciFact+AIGC dataset. The re-ranking methods rerank the top-100 retrieved hits from a first-stage BM25 model.
Metrics Target Corpus Llama2-generated ChatGPT-generated
BM25 +MiniLM +monoT5 BM25 +MiniLM +monoT5
NDCG@1 Human-Written 26.7 21.3 19.7 24.3 18.3 21.3
LLM-Generated 21.0 32.7 39.7 24.3 35.7 39.3
$\text{Relative}~{}~{}\Delta$ 23.9 -42.2 -67.3 0.0 -64.4 -59.4
NDCG@3 Human-Written 40.3 42.8 45.9 38.5 41.4 46.4
LLM-Generated 38.8 47.8 52.9 40.2 50.1 54.2
$\text{Relative}~{}~{}\Delta$ 3.8 -11.0 -14.2 -4.3 -19.0 -15.5
NDCG@5 Human-Written 44.4 46.9 49.0 42.7 45.6 48.9
LLM-Generated 41.5 50.2 54.7 42.7 53.0 56.1
$\text{Relative}~{}~{}\Delta$ 6.8 -6.8 -11.0 0.0 -15.0 -13.7
MAP@1 Human-Written 25.7 20.8 18.9 23.7 17.9 20.5
LLM-Generated 19.6 30.8 37.8 23.1 33.8 37.8
$\text{Relative}~{}~{}\Delta$ 26.9 -38.8 -66.7 2.6 -61.5 -59.3
MAP@3 Human-Written 36.7 37.5 39.7 34.8 35.8 40.3
LLM-Generated 34.3 43.6 48.9 35.8 45.9 50.0
$\text{Relative}~{}~{}\Delta$ 6.8 -15.0 -20.8 -2.8 -24.7 -21.5
MAP@5 Human-Written 39.1 40.0 41.6 37.3 38.3 41.7
LLM-Generated 35.9 45.0 50.1 37.3 47.6 51.4
$\text{Relative}~{}~{}\Delta$ 8.5 -11.8 -18.5 0.0 -21.7 -20.8
@@ 3.3. Bias in Re-Ranking Stage
In a typical IR system, there are two primary stages of document filtering. The first stage involves a retriever, responsible for document recall, while the second stage employs a re-ranker, which fine-tunes the ordering of documents within the initially retrieved set. While we have revealed the presence of the source bias in the first stage, a natural pivotal research question remains: does this bias also manifest in the re-ranking stage? To delve into this, we select two representative and state-or-the-art re-ranking models: MiniLM (Wang et al., 2020) and monoT5 (Nogueira et al., 2020) to rerank the top-100 document list retrieved by a first-stage BM25 model.
The results on the SciFact+AIGC dataset with Llama-generated corpus and ChatGPT-generated corpus are presented in Table 6. From the results, while even the first-stage retrievers (BM25) may exhibit a preference for human-written content, the second-stage re-rankers once again demonstrate a bias in favor of LLM-generated content. Remarkably, the bias in re-ranking models appears to be more severe, as evidenced by the relative percentage difference of $-67.3\%$ and $-59.4\%$ in NDCG@$1$ for monoT5, respectively. These findings further confirm the pervasiveness of source bias in neural ranking models that rely on PLMs, regardless of the first retrieval stage or second re-ranking stage.